# Query Rewriting via LLMs

### v2

@@ Abstract
We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from standard benchmarks on contemporary database platforms. The results show significant performance improvements for slow queries, with regard to both abstract costing and actual execution, over both SOTA techniques and the native query optimizer. For instance, with TPC-DS on PostgreSQL, the geometric mean of the runtime speedups for slow queries was as high as 18.4 over the native optimizer, whereas SOTA delivered 6 in comparison.
@@ 1 Introduction
As a case in point, consider the blog-processing query shown in Figure 2, which computes a daily summary of rating metrics for highly-rated blogs. When this seemingly complex query is given to the current PostgreSQL engine (v16) [1], its execution plan essentially mimics the hierarchical structure of the query. This strategy leads to multiple scans and joins of the Blogs and Posts tables, making the query take several minutes to complete. However, the query can be equivalently rewritten (assuming NOT NULL column constraints) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
@@
A variety of innovative Rule-based (e.g. Learned Rewrite [45]) and Model-based (e.g. Gen-Rewrite [25]) approaches have been proposed in the literature for S2S transformations. Further, a recent vision paper [46] has persuasively advocated the use of LLMs for such query rewriting. These novel state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting. However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space rather than directly in query space – the latter provides greater scope for candidate rewrites since only intent, and not implementation, is expressed. In this paper, we address these lacuna and amplify the effectiveness of query rewriting.
@@
We propose LITHE (LLM Infused Transformations of HEfty queries), an LLM-based query rewriting assistant to aid DBAs in tuning slow-running queries that have entailed their intervention. As illustrated in the architectural diagram of Figure 3, LITHE takes as input the user query $Q^{U}$ and outputs a transformed query $Q^{T}$, together with (a) the expected performance improvement, in terms of optimizer estimated cost, of $Q^{T}$; (b) a verification label indicating the mechanism (provable or statistical) used to determine that $Q^{T}$ is semantically equivalent to $Q^{U}$; and (c) a reasoning for why the LLM expects $Q^{T}$ to be helpful wrt performance. Armed with this information, DBAs can leverage their expertise to decide whether or not to use $Q^{T}$.
@@
In particular, we work with two classes of rules:
1. Redundancy Removal Rules: These rules are meant to eliminate repeated and redundant computations of the same output.
2. Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts. To our knowledge, such metadata inclusion in prompts has not been considered before in the SQL rewrite context. As shown later in our experiments, it proves to be a powerful mechanism for ensuring performant rewrites across database environments.
Token Probability Driven Rewrites. The above rewrite options are restricted to the standard prompt interface. Additionally, we leverage the rich telemetry provided by LLMs – in particular, the token probabilities output at each step in the prediction sequence. Whenever the LLM lacks high confidence in the next token, we follow multiple alternative paths in the decision process. To ensure practical overheads on this enumerative approach, a Monte Carlo tree search (MCTS) technique is incorporated in our implementation. MCTS has previously demonstrated significant effectiveness in code generation exercises [43].
Semantic Equivalence Labels. After generating a rewrite, it is imperative to check for semantic equivalence to the original query. Unfortunately, the general problem of proving query equivalence is NP-complete [6]. However, for restricted classes of queries, various logic-based equivalence-proving tools are available, and in LITHE, we use the recently developed QED tool [36] to verify rewrite correctness for queries within its scope.
The remaining out-of-scope queries are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases constructed using down-sampled versions of the original database. Finally, result equivalence can be optionally tested on the original database itself. While these checks do not suffer from false negatives, they may (rarely) incur false positives.
After carrying out the above exercise, LITHE provides a verification label – provable or statistical – regarding semantic equivalence. Ultimately, it is left to the DBA’s discretion to determine the validity and utility of the rewrite.
@@
We also evaluated whether the above cost reductions translated to real execution speedups. Here, we find that LITHE is indeed often substantively faster at run-time as well. Specifically, the geometric mean of the runtime speedups for slow queries was as high as 18.4 over the native optimizer, whereas SOTA delivered 6 in comparison. However, we also found that the well-known discrepancies between optimizer cost predictions and actual execution times [26] could be exacerbated by an injudicious choice of rule formulations. This separation suggests a future line of research to reduce the reasoning distance between language-based foundation models and statistics-based optimizer models.
@@ 2 LITHE Overview
3. Semantic Equivalence. A suite of provable (logic-based) and statistical (result-based) techniques are employed to assess the semantic equivalence of a recommended rewrite. If the rewrite is deemed valid by this module, it is returned along with the prompt that generated it; otherwise an invalid label is returned.
@@
5. Output. The least expensive valid rewrite as identified by the MCTS module is returned to the DBA together with the cost of the rewritten query and the approach used to claim equivalence. Additionally, an LLM-generated reasoning for the expected performance improvement is also provided. If no valid rewrite is identified, the original query itself is returned.
@@ 3 Basic Prompts
The performance of the four prompt templates on the micro-benchmark is shown in Table 1. We find that less than half the rewrites are productive with individual prompts. However, a drill-down shows that the best prompt in the ensemble is query-specific – this opens up the possibility of using all four prompts in parallel, and then choosing the best among them. This ensemble approach raises the CPRs to 6 (Row 5 in Table 1) – however, there remain four queries that are not productively rewritten by these prompts.
@@ 4.1 Redundancy Removal
A natural question here would be whether, while retaining the one-rule-per-prompt design, the rules could be progressively applied with the output of one prompt provided as input to the next, and so on. This approach would benefit queries with multiple types of redundancies. However, it also increases rewrite overheads due to repeated prompt applications. So, for simplicity, we have chosen to process them individually rather than cumulatively.
@@ 4.2 Selectivity-based Guidance
    Selectivities of columns appearing in WHERE and JOIN clauses, as obtained from the database engine.
@@
We note in closing that rules R1 through R6 not only add queries to the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3. Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule – this option is examined in the Section B.
@@ 5 Token Probability Driven Rewrite
A key challenge when using LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up. This is often due to the output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs. In order to have a robust approach for such cases, we take inspiration from the code generation literature [43]. Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite as well as performance improvements.
@@
11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n,a)$
@@ 5.1 MCTS Search Process
3. Simulation: Here, we determine the potential value, $V(n_{cur})$, to be assigned to $n_{cur}$. This node is expanded in a greedy fashion, based on the highest-probability tokens until a terminal node is reached (Line 21). Then, the complete rewritten query represented by the terminal node’s state is used to compute the potential (Lines 22–23). For a valid rewrite, $V(n_{cur})$ is equal to the speedup it provides with respect to the original query. However, if invalid (i.e. syntactically or semantically incorrect), $V(n_{cur})$ is assigned a zero value.
After every simulation, the complete rewritten query obtained after the greedy expansion of $n_{cur}$ is cached along with $V(n_{cur})$ in a map, $Potential$.
4. Back Propagation: The $V$ value of the simulation for $n_{cur}$ is backpropagated to all its ancestor nodes. An ancestor node’s $V$ value is updated if and only if the new value is higher than the existing value.
@@ 5.3 Performance
We find that LITHE’s performance with MCTS-based rewrites fully matches the human target, providing 10 CPRs and the maximum CSGM of 11.84. While these improvements, as compared to just prompting, might seem marginal at this stage, we wish to make two observations: (1) MCTS helps extract the maximum improvement wherever possible; and (2) As shown later in Section 7.5.2, improvements due to MCTS become significant for smaller LLMs such as LLaMA.
@@ 6.1 LITHE Parameter Settings
The “temperature” parameter of GPT-4o, which ranges over [0,1], controls the randomness of the model’s response. While a higher temperature can be useful for creative writing where one would seek diverse and exploratory answers, in our case we want a focused and deterministic answer as far as possible. Hence we set the temperature to 0 which forces the model to greedily sample the next token.
@@ 6.2 Query Equivalence Testing
We use a multi-stage approach, described below, to test semantic equivalence between the original query and a candidate rewrite.
1. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is NP-complete [6], a variety of logic-based tools (e.g. Cosette[11], SQL-Solver [15], VeriEQL [17], QED [36]) are available for proving equivalence over restricted classes of queries, as mentioned in the Introduction. In LITHE, we use the recently proposed QED [36] since it was found to cover a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome and relatively inexpensive.
2. Result Equivalence via Sampling. If the original query is not within the QED scope, we alternatively use a sampling-based approach to test equivalence. The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not sufficient. That is, false positives may be present because the sampled database may not cover all the predicates featured in the query. To minimize this likelihood, we use a combination of (1) correlated sampling [42] for maintaining join integrity in the sample, (2) adding synthetic tuples in the sample to distinguish outer and inner joins, and (3) adjusting constants in the filter predicates to produce populated results – the complete details are in the Section C.
3. Result Equivalence on the Entire Database. Result equivalence could also be evaluated, in principle, on the entire database itself. Of course, this could turn out to be prohibitively expensive, especially if the queries themselves are time-consuming (e.g. due to the scale of the underlying database) and/or if the candidate rewrites happen to be regressions. Therefore, we leave this check as an optional choice for the DBA.
@@ 7 Experimental Evaluation
LITHE has also been evaluated on other benchmarks, including DSB [13], ARCHER [44], JOB [19] and StackOverflow [27]. Due to space limitations, we defer their details to Section A.
@@
Metrics. For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with ¿ 1.5 speedup) could be constructed. Subsequently, we computed the CSGM (Cost Speedup Geometric Mean) and TSGM (Time Speedup Geometric Mean) performance obtained by each technique over the set of all CPRs (i.e. CPRs arising from either LITHE or SOTA).
@@ 7.1.1 Estimated Cost
LITHE produces a rewrite with a positive speedup ($>$ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold. Of these 46, there were 26 CPRs resulting in a highly productive CSGM of 11.5. On the other hand, SOTA delivers only 13 CPRs (out of 42 positive rewrites) with a CSGM of 6.1. All but one of the SOTA CPRs also feature in the LITHE CPRs, making the total number of CPRs considered to be 27.
We were able to prove the rewriting correctness of 9 CPRs through the logic-based QED tool. The remaining passed our result equivalence tests on both the down-sampled and full databases. Furthermore, as a final check, we manually verified the correctness of the rewritten queries.
@@
Second, in most queries, the cost speedup of LITHE either exceeds or matches SOTA. In fact, there are several queries (e.g. Q45, Q25, Q4) where LITHE produces a highly beneficial CPR but SOTA simply returns the original query. Conversely, the opposite is true for one query (Q57) where SOTA projects a large speedup but LITHE settles for the original query. And in Q88 and Q95, SOTA is marginally better than LITHE.
At this stage, one could imagine that more rules could be added to LITHE to bring it on par with SOTA for queries like Q57. However, as discussed later in Section 8.3, such customized rules, while attractive at compile-time, run a significant risk of causing regressions at run-time. We have therefore selected only broad-brush rules for inclusion in LITHE. This conservative approach is justified below in the timing section, where we find that LITHE actually performed better than SOTA on these queries (Q57, Q88, Q95) wrt runtimes.

### v3

@@ Abstract
We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from standard benchmarks on contemporary database platforms. The results show significant performance improvements for slow queries, with regard to both abstract costing and actual execution, over both SOTA techniques and the native query optimizer. For instance, with TPC-DS on PostgreSQL, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison.
@@ 1 Introduction
As a case in point, consider the blog-processing query shown in Figure 2, which computes a daily summary of rating metrics for highly-rated blogs. When this seemingly complex query is given to the current PostgreSQL engine (v16) [2], its execution plan essentially mimics the hierarchical structure of the query. This strategy leads to multiple scans and joins of the Blogs and Posts tables, making the query take several minutes to complete. However, the query can be equivalently rewritten (assuming NOT NULL column constraints and key-joins) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
@@
A range of innovative Rule-based (e.g. Learned Rewrite [48]) and Model-based (e.g. Gen-Rewrite [28]) approaches have been proposed for S2S transformations. A recent vision paper [49] also advocates the use of LLMs for such query rewriting. The state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting. However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space (i.e optimization on the nodes of the execution plan tree) rather than directly in query space (i.e transforming the query structure itself). In this paper, we address these lacuna and amplify the effectiveness of query rewriting. In particular, our query space-based transformations provide greater scope for candidate rewrites since only intent, and not implementation, is expressed, and this multiplicity of rewrites in turn results in a broader coverage of the underlying plan space.
@@
We propose LITHE (LLM Infused Transformations of HEfty queries), an LLM-based query rewriting assistant to aid DBAs in tuning slow-running queries that have entailed their intervention. As illustrated in the architectural diagram of Figure 3, LITHE takes as input the user query $Q^{U}$ and outputs a transformed query $Q^{T}$, together with (a) the expected performance improvement, in terms of optimizer estimated cost, of $Q^{T}$; (b) a verification label indicating the mechanism (provable or statistical) used to determine that $Q^{T}$ is semantically equivalent to $Q^{U}$; and (c) a reasoning for why the LLM expects $Q^{T}$ to be helpful wrt performance. Armed with this information, DBAs can leverage their expertise to decide whether or not to use $Q^{T}$. Note that having the DBA in the loop is a common practice in commercial query advisory systems [14].
@@
In particular, we work with two classes of rules – 1. Redundancy Removal Rules: These rules eliminate repeated and redundant computations of the same output; and 2. Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts. To our knowledge, such metadata inclusion in prompts has not been considered before in the SQL rewrite context. As shown later in our experiments, it proves to be a powerful mechanism for ensuring performant rewrites across database environments.
Token Probability Driven Rewrites. The above rewrite options are restricted to the standard prompt interface. We also leverage the rich telemetry provided by LLMs – particularly, the token probabilities output at each step in the prediction sequence. Whenever the LLM lacks high confidence in the next token, we follow multiple alternative paths in the decision process. To ensure practical overheads on this enumerative approach, a Monte Carlo tree search (MCTS) technique is incorporated in our implementation. MCTS has previously been used to improve ordering of query rewrite rules [48] in the plan space, but our usage is new – it is geared to guide the LLM in exploring rewrites in the query space.
Semantic Equivalence Labels. After generating a rewrite, it is imperative to check for semantic equivalence to the original query. Unfortunately, the general problem of proving query equivalence is NP-complete [7]. We therefore take a twofold approach to check for query equivalence. First, the queries are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases constructed using down-sampled versions of the original database. While these checks do not suffer from false negatives, they may (rarely) incur false positives. These tests can thus quickly weed out obviously non-equivalent rewrites.
A final check is then performed using logic-based equivalence-proving tools that provably work for a restricted classes of queries. Specifically, LITHE uses the recently developed QED tool [40] and SQLSolver [17] to verify rewrite correctness for queries within their scope. Whereas, for out-of-scope queries, result equivalence on the original database itself is used to test the equivalence.
After carrying out the above exercise, LITHE provides a verification label of semantic equivalence – provable if verified by a logic-based tool, or statistical otherwise. Ultimately, it is left to the DBA’s discretion to determine the validity and utility of the rewrite.
Regression Identification. Finally, it is well-known that in practice there are often material discrepancies between optimizer cost predictions and actual execution times [29]. So it is possible that a promising rewrite as per the optimizer may actually regress at run-time as compared to the original formulation. To help guard against such possibilities, we have implemented heuristic mechanisms, described in Section 6.3, to identify “brittle” rewrites and discard them in favor of the original.
@@
We also evaluated whether the above cost reductions translated into real execution speedups. Here, we find that LITHE is indeed often substantively faster at run-time as well. Specifically, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison. Finally, our regression identification techniques were successful in guarding against “brittle” rewrites that are promising cost-wise, but eventually poor at run-time.
@@ 2 LITHE Overview
3. Fast Semantic Equivalence. Statistical (result-based) techniques are employed to quickly and cheaply assess the semantic equivalence of a recommended rewrite. If the rewrite is deemed valid by this module, it is returned along with the prompt that generated it; otherwise an invalid label is returned.
@@
5. Final Checks and Output. The least expensive valid rewrite (as identified by the MCTS module) is tested using a suite of provable (logic-based) techniques to generate a final equivalence label (provable or statistical), and its cost benefit over the original query is computed. Further, the execution “brittleness” is assessed using the robustness heuristics. If no valid rewrite is identified, or if the rewrite is expected to be a regression, the original query itself is returned to the DBA. Whereas, if a beneficial rewrite is recommended, an LLM-generated reasoning for the expected performance improvement is also extracted.
@@ 3 Basic Prompts
Table 1 shows the performance of the four prompt templates on the micro-benchmark. We find that less than half the rewrites are productive with individual prompts. However, a drill-down shows that the best prompt in the ensemble is query-specific – this opens up the possibility of using all four prompts in parallel, and then choosing the best among them. This ensemble approach raises the CPRs to 6 (Row 5 in Table 1) – however, there remain four queries that are not productively rewritten by these prompts.
@@ 4.1 Redundancy Removal
A natural question here would be whether, while retaining the one-rule-per-prompt design, the rules could be progressively applied with the output of one prompt provided as input to the next, and so on. This approach would benefit queries with multiple types of redundancies. However, it also introduces significant computational overheads due to the vast number of possible rule application orders, making it expensive to explore. So, for simplicity, we have chosen to process them individually rather than cumulatively.
@@ 4.2 Selectivity-based Guidance
    Estimated selectivities of columns appearing in the WHERE and JOIN clauses – these values are obtained via calls to the cardinality estimation modules of the query optimizer.
@@
We note in closing that rules R1 through R6 not only add queries to the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3. Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule – this option is examined in Section B.
@@ 5 Token Probability Driven Rewrite
A key challenge with LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up. This is often due to the output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs [20]. In order to have a robust approach for such cases, we take inspiration from the code generation literature [46]. Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite as well as performance improvements. As mentioned earlier, MCTS has been previously used for ordering rule applications in query rewrites [48], whereas our goal is to guide the LLM in exploring query rewrites.
@@
11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$
@@ 5.1 MCTS Search Process
3. Simulation: Here, we determine the potential value, $V(n_{cur})$, to be assigned to $n_{cur}$. This node is expanded greedily based on the highest-probability tokens until a terminal node is reached (Line 21). Then, the complete rewritten query represented by the terminal node’s state is used to compute the potential (Lines 22–23). For a valid rewrite, $V(n_{cur})$ is equal to the speedup it provides w.r.t the original query. However, if invalid (i.e. syntactically or semantically incorrect), $V(n_{cur})$ is assigned a zero value. After every simulation, the complete rewritten query obtained after the greedy expansion of $n_{cur}$ is cached along with $V(n_{cur})$ in a map, $Potential$.
4. Back Propagation: The $V$ value of the simulation for $n_{cur}$ is back-propagated to all its ancestor nodes, and their values are updated iff the new value is higher than the existing value.
@@ 5.3 Performance
LITHE’s performance with MCTS-based rewrites fully matches the human target, yielding 10 CPRs and the maximum CSGM of 11.84. While these gains over just prompting may seem marginal, we wish make two observations: (1) MCTS extracts maximum improvement whererever possible, and (2) As shown in Section7.5.2, its impact becomes significant for smaller LLMs like LLaMA.
@@ 6.1 LITHE Parameter Settings
The “temperature” parameter of GPT-4o, which ranges over [0,1], controls randomness of the model’s response. While higher values suit creative tasks that require diverse and exploratory outputs, we seek focused, deterministic answers. Thus, we set the temperature to 0, forcing the model to greedily sample the next token.
@@ 6.2 Query Equivalence Testing
We use a multi-stage approach, described below, to help the DBA test semantic equivalence between the original query and a recommended rewrite. Note that, as mentioned earlier, having the DBA in the loop is a common practice in commercial query advisory systems[14].
1. Result Equivalence via Sampling. We use a sampling-based approach to quickly test equivalence in the rewrite generation stages of the pipeline. The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not sufficient. That is, false positives may be present because the sampled database may not cover all the predicates featured in the query. To minimize this likelihood, we use a combination of (1) correlated sampling [45] for maintaining join integrity in the sample, (2) adjusting constants in the filter predicates to produce populated results, and (3) injecting query-specific synthetic tuples in the sample, similar to the XData mutant-killing tool [34], to cover predicate boundary conditions – the complete details are in Section C.
2. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is NP-complete [7], a variety of logic-based tools (e.g. Cosette[12], SQL-Solver [17], VeriEQL [21], QED [40]) are available for proving equivalence over restricted classes of queries. Once the least-cost rewrite is identified, LITHE uses the recently proposed QED [40] and SQLSolver [17] since these two approaches were found to cover a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome. Note that this rewrite has already passed the sampling-based tests described above.
3. Result Equivalence on the Entire Database. If the logic-based test is inconclusive, result equivalence is evaluated on the entire database itself. The DBA may choose to prematurely terminate this test in case the checking time is found to be excessive.
6.3 Regression Identification
We use heuristics to predict whether the actual execution time of a given promising rewrite may turn out to be slower than the original query. The first heuristic is based on the SEER robust plan identification algorithm [19], which is predicated on the database engine providing a “force-plan” feature (e.g. [1]). Let $P^{U}$ and $P^{T}$ be the plans corresponding to the original query, $Q^{U}$, and the recommended rewrite, $Q^{T}$, respectively. We construct parametrized versions of these queries, where the constants in the filter predicates are replaced by variables. Then, by assigning appropriate values to these parameters, we construct queries that are located at the corners of the selectivity space. The plans $P^{U}$ and $P^{T}$ are forced at each of these corner locations, and if the rewrite’s cost is lower than the original at all of them, the rewrite is deemed to be robust.
At this time, plan forcing is not available in all database engines. For such limited environments, we fall back to an alternative heuristic – we compare the query runtimes obtained on the series of executions carried out on sampled databases during semantic equivalence testing. If the rewrite’s runtimes are lower in all of these intermediate executions, then it is estimated to be robust.
@@ 7 Experimental Evaluation
LITHE has also been evaluated on other benchmarks, including DSB [15], ARCHER [47], JOB [23] and StackOverflow [30]. The overall performance characteristics were found to be similar to those presented here – due to space limitations, we defer their details to Section A.
@@
Metrics. For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with $>$ 1.5 speedup) could be constructed. Subsequently, we computed the CSGM (Cost Speedup Geometric Mean) and TSGM (Time Speedup Geometric Mean) performance obtained by each technique over the set of all CPRs (i.e. CPRs arising from either LITHE or SOTA).
@@ 7.1.1 Estimated Cost
LITHE produces a rewrite with a positive cost speedup ($>$ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold. Of these 46, there were 26 CPRs resulting in a highly productive CSGM of 11.5. On the other hand, SOTA delivers only 13 CPRs (out of 42 positive rewrites) with a CSGM of 6.1. All but one of the SOTA CPRs also feature in the LITHE CPRs, making the total number of CPRs considered to be 27. Of these 27, we were able to formally verify 11 using the logic-based tools, whereas the remaining 16 passed our statistical tests. Furthermore, we also manually verified the correctness of these rewritten queries.
@@
Second, in most queries, LITHE’s cost speedup either exceeds or matches SOTA. In fact, for several queries (e.g. Q45, Q25, Q4), LITHE produces a highly beneficial CPR but SOTA returns the original query. Conversely, the opposite is true for one query (Q57) where SOTA projects a large speedup but LITHE settles for the original query. And in Q88 and Q95, SOTA performs only marginally better.
At this stage, one might expect that adding more rules to LITHE could bring it on par with SOTA for queries like Q57. However, we deliberately include only broad-brush rules in LITHE to ensure generalizability and efficiency. As the following timing section shows, due to this conservative approach LITHE actually outperforms SOTA on these queries (Q57, Q88, Q95) in terms of runtime.